OpenStudy (anonymous):

Find the derivative of (sin x)^1/x.

OpenStudy (anonymous):

That's the function you want to derivate??\[(\frac{ \sin x }{ x})'\]

OpenStudy (anonymous):

the equation is\[(\sin x)^\frac{ 1 }{ x }\]

OpenStudy (dumbcow):

chain rule let u = sin x ..... du = cos x \[\frac{d}{dx} u^{1/x} = \frac{1}{x}u^{1/x - 1} * du\]

OpenStudy (dumbcow):

wait i think i messed that up :{

OpenStudy (anonymous):

so would the answer be: \[\frac{ 1 }{ x }(\sin x)^{\frac{ 1 }{ x }-1}\cos x\] ?

OpenStudy (dumbcow):

yeah but i dont think i did derivative right for u^1/x

OpenStudy (anonymous):

yeah, that's the part that fried my brain :/ not sure how to approach the 1/x

OpenStudy (dumbcow):

ok i got it, its just another level of chain rule let w = 1/x ..... dw = -1/x^2 \[\frac{d}{dx} u^w = \ln (u) u^w * du * dw\]

OpenStudy (anonymous):

ahhh.. I see it! Thank you so much!!

OpenStudy (dumbcow):

your welcome

OpenStudy (alekos):

where did the ln(u) come from?

OpenStudy (dumbcow):

derivative of an exponential function \[\frac{d}{dx} a^x = \ln(a) a^x\] thats why derivative of e^x is just e^x

OpenStudy (alekos):

yes, of course

OpenStudy (alekos):

that was well thought out. who comes up with these problems?

OpenStudy (anonymous):

Monash University :( I have a maths assignment.. My heart broke when I saw the questions

OpenStudy (alekos):

ah! that explains it. They dont make things easy

OpenStudy (anonymous):

not at all